I needhelp setting up this word problem.
You have to sell 50 tickets to a fair. Student tickets cost $3 and adult tickets cost $5. Let t represent the number of student tickets you sell. Write and simplify an expression in terms of t for the total amount of money you raise for selling 50 tickets. Then use the expression to determine how much money you raise if you sell all 50 tickets and 28 of them are student tickets.
I have figured out that if I sell 28 student tickets I will sell 22 adult tickets and I will raise $194. 50-28=22 28(3) + 22(5) = 194
I just can't figure out what the first equation is.

Answer by stanbon(75887)

I needhelp setting up this word problem.
You have to sell 50 tickets to a fair. Student tickets cost $3 and adult tickets cost $5. Let t represent the number of student tickets you sell. Write and simplify an expression in terms of t for the total amount of money you raise for selling 50 tickets. Then use the expression to determine how much money you raise if you sell all 50 tickets and 28 of them are student tickets.
I have figured out that if I sell 28 student tickets I will sell 22 adult tickets and I will raise $194. 50-28=22 28(3) + 22(5) = 194
----------------
Quantity Equation:
t + a = 50
----
If t = 28, a = 22
Value is 3*28 + 5*22 = $194
----
Cheers,
Stan H.
----